Your agency is paid on a traditional 15 percent media commission. You purchase $300,000 in network TV time to run an ad for your client. How much will your client pay the agency for the TV time? How much will the network receive? How much will the agency keep?The ad you ran cost $30,000 to produce. Your agency charges a 17.65 percent markup to the client for outside production costs. How much will your client pay the agency for the production? How much of that will your agency keep?

1
Expert's answer
2021-05-25T03:00:02-0400

Cost of running an ad for the client in TV by the agency = $300,000

Cost to client will be marked up by 15% "=300,000*1.15 = 345,000"

Agency will receive $345,000

Network TV will receive $300,000

The agency will keep = "345,000-300,000 = 45,000"

Hence $45,000 will be the profit to the agency.

Cost of producing the ad = $30,000

Charge to the client "=30,000*1.1765 = 35,295"

Client will keep = "35,295-30,000"

Therefore the client will make profit of $5,295